# Comparison of chiral limit studies in curvature mass versus on-shell renormalized quark-meson model using ChPT

### v1

@@ II.3 The ChPT scaling of $\bf f_{\pi},f_{K}\ \text{and}\ M_{\eta}^{2}=m_{\eta}^{2}+m_{\eta^{\prime}}^{2}$
In our very recent 2+1 flavor RQM model work vkt25 , we investigated the chiral limit and generated the $m_{\pi}-m_{K}$ ( $m_{ud}-m_{s}$) planes of the Columbia plot where two input sets of the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results from the infrared regularized U(3) ChPT and the standard large $N_{c}$ U(3) ChPT, were used for obtaining the ($m_{\pi},m_{K}$) dependence of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ when one moves away from the physical point. Though the $\eta^{\prime}$ meson is treated differently in the two ChPT approaches, only marginal differences were found in the extent of first, second order and crossover transition regions of the Columbia plot when the RQM model results from two input sets were compared in Ref. vkt25 . Therefore in the present QMVT model study, the chiral limit and the $m_{\pi}-m_{K}$ ($m_{ud}-m_{s}$) planes of the Columbia plot, will be investigated using only the infrared regularized U(3) ChPT given ($m_{\pi},m_{K}$) dependent expressions of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ which are briefly reproduced below.
@@
In the above, the chiral logarithms $\mu_{\text{\tiny{PS}}}=\frac{m_{\text{\tiny{PS}}}^{2}}{32\pi^{2}}\ln(\frac{m_{\text{\tiny{PS}}}^{2}}{M_{0}^{2}})$ are at the scale $M_{0}$ and for the $m_{\text{\tiny{PS}}}^{2}$, one takes the leading order squared mass of the corresponding meson of the pseudo-scalar octet. Recall that the chiral constants $L_{i}$ do not vary with the pseudo-scalar masses. The parameters $q=2m_{s}/(m_{u}+m_{d})=m_{s}/m_{ud}$ and $A=B\ (m_{u}+m_{d})/2=B\ m_{ud}$ where the constant $B$ is determined by the quark condensate $<\bar{u}u>$ in the chiral limit, govern the quark mass and quark condensate dependencies of the above one-loop ChPT expressions of the $m_{\pi}^{2},m_{K}^{2},f_{\pi}\ \text{and}\ f_{K}$. The $q$ and $A$ can be expressed below in terms of the masses $m_{\pi},m_{K},m_{\eta}$ and chiral constants $L_{i}$ after inverting the Eqs. (24) and (25) to order $\mathcal{O}(\frac{1}{f^{2}})$ accuracy.
@@
With the input $f_{\pi}$=93 MeV, $f_{K}$=113 MeV, $m_{\pi}$=138 MeV, $m_{K}$=495.6 MeV,$m_{\eta}$=547.8 MeV and $M_{0}=4\pi f_{\pi}\equiv 1168$ MeV, $f$=88 MeV, one gets the constants $L_{4}$ and $L_{5}$ at the physical point as :
@@
$\frac{1}{f^{2}}\biggl{[}8v_{0}^{(2)}(2m_{K}^{2}+m_{\pi}^{2})(3L_{4}+L_{5})+m_{\pi}^{2}(\mu_{\eta}-3\mu_{\pi})$
$-4m_{K}^{2}\mu_{\eta}+\frac{16}{3}\biggl{\{}(6L_{8}-3L_{5}+8L_{7})(m_{\pi}^{2}-m_{K}^{2})^{2}+$
$2L_{6}(m_{\pi}^{4}-2m_{K}^{4}+m_{\pi}^{2}m_{K}^{2})+L_{7}(m_{\pi}^{2}+2m_{K}^{2})^{2}\biggr{\}}\biggr{]}\;.$
The $\eta$ mass occurs at the leading order in chiral logarithm i.e. $(m_{\eta}^{(0)})^{2}=(4m_{K}^{2}-m_{\pi}^{2})/3$, hence the functions $f_{\pi}(m_{\pi},m_{K})$, $f_{K}(m_{\pi},m_{K})$ and $M_{\eta}^{2}(m_{\pi},m_{K})$ are applicable only when $4m_{K}^{2}>m_{\pi}^{2}$ The Eqs.(30) and (31) together with the Eqs.(35), enable the calculation of the QM/QMVT/RQM model couplings $\lambda_{2},c$,$m^{2}$ and $\lambda_{1}$ in the $m_{\pi}-m_{K}$ plane.
@@
[…]
@@
III Results and Discussions
Several QMVT model studies vac ; lars ; schafwag12 ; chatmoh1 ; guptiw ; vkkr12 ; vkkt13 ; Rai , where the QM model parameters are fixed using the curvature masses of the mesons in the presence of fermionic vacuum fluctuations, have shown that the quark one-loop vacuum fluctuations cause very large smoothing effect on the temperature and chemical potential variations of the non-strange, strange condensates, their derivatives and the mass $m_{\sigma}$ of the scalar $\sigma$ meson. With the parameters obtained using the physical point $\pi,K,\eta\text{ and }\eta^{\prime}$ masses, the 2+1 flavor QMVT and RQM model order parameters, its derivatives and phase diagrams in the $\mu-T$ plane, have been compared extensively in Ref.vkkr22 ; skrvkt24 and it has been shown that the critical end point (where the crossover transition turns second order), shifts to very large chemical potentials and smaller temperatures in the curvature mass parametrized QMVT model where the chiral crossover transition on the temperature axis at $\mu=0$, occurs at a noticeably larger pseudo-critical temperature $T_{c}^{\chi}$ in comparison to the RQM model. Recall that the smoothing effect of the quark one-loop vacuum fluctuation, on the chiral transition is moderate in the RQM model skrvkt24 because on-shell method of renormalizing the parameters in the RQM model, generates a significantly stronger axial $U_{A}(1)$ anomaly t’Hooft coupling $c$ while the explicit chiral symmetry breaking strength $h_{x}$ ($h_{y}$) in the non-strange (strange) direction becomes weaker by a small amount (a relatively large amount). Note that the $c,h_{x}\text{ and }h_{y}$ do not change in the QMVT model at physical point from their tree level values obtained in the QM model.
It is worthwhile to perform the chiral limit study and draw the Columbia plot in the curvature mass parametrized QMVT model where the infrared regularized U(3) ChPT is employed to determine the ($m_{\pi},m_{K}$) dependent numerical values of the $f_{\pi},f_{K}\text{ and }M_{\eta}^{2}$ when one reduces the $m_{\pi}$ and $m_{K}$ while moving away from the physical point. The softening effect of the fermionic vacuum fluctuations in the QMVT model needs to be numerically quantified by calculating its effect on the location of the tricritical point (TCP) at $\mu=0$ in the light chiral limit and the value of the pion (light quark) critical mass $m_{\pi}^{c}$ ($m_{ud}^{c}$) beyond which the chiral transition is a smooth crossover on the three flavor chiral limit line ($m_{\pi}=m_{K}$ i.e. $m_{ud}=m_{s}$ ) in the Columbia plot. In order to see how the chiral limit results are qualitatively and quantitatively influenced by the different methods of treating the quark one-loop vacuum fluctuation when it gets incorporated in the QM Model, the $\mu-m_{K}$, $m_{\pi}-m_{K}$ and the light-strange quark mass $\mu-m_{s}$, $m_{ud}-m_{s}$ planes of the Columbia plots, the $m_{K}^{TCP}(m_{s}^{TCP})$ location of the TCP on the $m_{K}(m_{s})$ axis where the $m_{\pi}(m_{s})=0$ and the $m_{\pi}^{c}$ ($m_{ud}^{c}$) obtained for the QMVT model calculation in this work will be compared with the corresponding results of the e-MFA:QM model FRG study in the Ref. Resch and the very recent RQM model chiral limit study reported in the Ref. vkt25 .
@@ III.1 QMVT versus RQM model order parameters and its derivatives in the path to chiral limit
The light chiral limit in the QMVT (RQM) model is given by $h_{x}=0$ ($h_{x0}=0$) while the strange chiral limit is given by $h_{y}=0$ ($h_{y0}=0$). The reduced chiral symmetry breaking strengths $h_{x}^{*}$ and $h_{y}^{*}$ ($h_{x0}^{*}$ and $h_{y0}^{*}$) and the corresponding QMVT (RQM) model parameters $\lambda_{2},c,\lambda_{1},m^{2}\ (\lambda_{20},c_{0},\lambda_{10},m_{0}^{2})$ for the chiral limit study, are calculated by using the infrared regularized U(3) ChPT described in the section II.3. The changes in pion, kaon decay constants $f_{\pi}$ and $f_{K}$ are calculated by putting the values of the reduced pion, kaon masses $m_{\pi}^{*}$ and $m_{K}^{*}$ respectively into the Eq. (30) and Eq. (31) while $\text{M}_{\eta}$ is calculated using the Eq.(35). The ratio $\frac{m_{\pi}^{*}}{m_{\pi}}=\frac{m_{K}^{*}}{m_{K}}=\beta$ is kept fixed while the starred $\pi$ and $K$ masses are reduced. The above choice gives us a direct path from the physical point to the chiral limit by just tuning only the fraction $\beta$ instead of randomly changing both the $m_{\pi}$ and $m_{K}$ and ensures that the $\frac{m_{\pi}^{*}}{m_{K}^{*}}$ remains equal to the ratio $\frac{m_{\pi}}{m_{K}}$ Schaefer:09 at the physical point where the ChPT parameter $q=\frac{2\ m_{s}}{m_{u}+m_{d}}=\frac{m_{s}}{m_{ud}}$ as the ratio of the strange $m_{s}$ to the average of light quark mass $m_{ud}$, also gets fixed. The above defined fraction $\beta$, also facilitates the comparison of our results with those of the Refs. Resch ; Schaefer:09 for the analogous fractions.
@@
The QMVT and RQM model temperature variations of the light and strange condensate $x$ and $y$ have been been compared respectively in the Fig. 1 and Fig. 1 when $\beta=1.0,0.5,0.3642,0.2028\text{ and }0.00725$ for the $m_{\sigma}=530$ MeV. When ChPT expression of $M_{\eta}$ is used as input to determine parameters of the QMVT and RQM models, the $x$ ($y$) plot at the physical point $\beta=1$, overlaps (not shown here), as reported in a very recent RQM model work in the Ref. vkt25 , with the corresponding temperature plot obtained after taking the $M_{\eta}:\text{Expt}$ in input for finding the model parameters whereas the $f_{\pi},f_{K}$ are the same as in the first row of Tables 1 and 2. The light and strange condensate temperature variations in the Fig. 1 and Fig. 1, show a very smooth and delayed melting in the QMVT model for the $\beta=1.0\text{ and }0.5$ while the corresponding melting in the RQM model is sharper and occurs early on the temperature axis. This is evident from the temperature variation of the derivative $\partial x/\partial T$ in the Fig. 2 where the chiral crossover transition in the QMVT model occurs at the pseudo-critical temperature (defined as the peak of the $\partial x/\partial T$ temperature variation) of $T_{c}^{\chi}=162.16$ MeV for $\beta=1$ and $T_{c}^{\chi}=158.32$ MeV for $\beta=0.5$ while a relatively sharper chiral crossover transition, with a higher peak in the $\partial x/\partial T$ temperature plot, occurs in the RQM model at about 23-24 MeV smaller transition temperature of $T_{c}^{\chi}=139.42$ MeV for $\beta=1$ and $T_{c}^{\chi}=133.97$ MeV for $\beta=0.5$ The $\partial x/\partial T$ peak height in the QMVT model respectively for the $\beta=0.4$ and $\beta=0.3$ in the Fig. 2 turn out to be about five and three times smaller than the height of the $\partial x/\partial T$ peak in the RQM model for the case of $\beta=0.4$ The reason of the above pattern lies in the fact that the $\beta=0.4$ is close to the RQM model $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{ RQM})}=0.3642\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(50.26,~{}180.64)$ MeV where the peak in the $\partial x/\partial T$ temperature plot of the Fig. 2, diverges at the $T_{c}^{\chi}=129.33$ MeV and the crossover transition ends at the second order $Z_{2}$ critical point to become first order transition for the $\beta<\beta_{cep}$. The RQM model $\partial x/\partial T$ plots for the $\beta=0.3748$ and $\beta=0.3816$ close to the $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{ RQM})}=0.3642$ are also shown in the Fig. 2 to have a perspective of the divergence. The fraction $\beta$ needs significant reduction to get the $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{QMVT})}=0.2028\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(27.99,~{}100.59)$ MeV in the QMVT model when the peak height of the $\partial x/\partial T$ plot diverges at larger $T_{c}^{\chi}=145.06$ MeV in the Fig. 2 where the adjacent QMVT model $\partial x/\partial T$ plots are shown for the $\beta=0.2131$ and $\beta=0.2200$ also. The Fig. 1 and 1 show the light and strange order parameter temperature variations also for the $\beta_{cep}$ and $\beta=0.00725$ where the order parameters have a very large first order gap when one reaches very close to the chiral limit. Note that for the smaller $\beta<\beta_{cep}$, the difference in the QMVT and RQM model critical temperatures for the chiral transition decreases from about 24 MeV for the $\beta=1.0\text{ and }0.5$ to 16.35 MeV for $\beta=0.15$ and 13.35 MeV for $\beta=0.00725$. In the exact chiral limit $f_{\pi}=f_{K}$, the $m_{\eta}=0$ and the heavy $m_{\eta^{\prime}}\equiv 825$ MeV signifies the $U_{A}(1)$ anomaly for both the QMVT model and RQM model. The problem of the loss of the spontaneous chiral symmetry breaking ( SCSB) when the $m_{\pi},m_{K}\rightarrow 0$ in the chiral limit for the moderate $m_{\sigma}=400-800$ MeV, gets cured after using the infrared regularized U(3) ChPT scaling Herpay:05 of the $m_{\pi},m_{K},M_{\eta}^{2}$.

### v2

@@ II.3 The ChPT scaling of $\bf f_{\pi},f_{K}\ \text{and}\ M_{\eta}^{2}=m_{\eta}^{2}+m_{\eta^{\prime}}^{2}$
In our very recent 2+1 flavor RQM model work vkt25 , we investigated the chiral limit and generated the $m_{\pi}-m_{K}$ ( $m_{ud}-m_{s}$) planes of the Columbia plot where two input sets of the $\mathcal{O}(\frac{1}{f^{2}})$ accurate results from the infrared regularized U(3) ChPT termed as the input set $\text{M}_{\eta}$-I in Ref.vkt25 and the standard large $N_{c}$ U(3) ChPT termed as the input set $\text{M}_{\eta}$-II in Ref.vkt25 , were used for obtaining the ($m_{\pi},m_{K}$) dependence of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ when one moves away from the physical point. Though the $\eta^{\prime}$ meson is treated differently in the two ChPT approaches, only marginal differences were found in the extent of first, second order and crossover transition regions of the Columbia plot when the RQM model results from two input sets were compared in Ref. vkt25 . Therefore in the present QMVT model study, the chiral limit and the $m_{\pi}-m_{K}$ ($m_{ud}-m_{s}$) planes of the Columbia plot, will be investigated using only the infrared regularized U(3) ChPT given ($m_{\pi},m_{K}$) dependent expressions of the $f_{\pi},f_{K}\text{ and }M_{\eta}$ which are briefly reproduced below.
@@
In the above, the chiral logarithms $\mu_{\text{\tiny{PS}}}=\frac{m_{\text{\tiny{PS}}}^{2}}{32\pi^{2}}\ln(\frac{m_{\text{\tiny{PS}}}^{2}}{M_{0}^{2}})$ are at the scale $M_{0}$ and for the $m_{\text{\tiny{PS}}}^{2}$, one takes the leading order squared mass of the corresponding meson of the pseudo-scalar octet. Recall that the chiral constants $L_{i}$ do not vary with the pseudo-scalar masses. The parameters $q=2m_{s}/(m_{u}+m_{d})=m_{s}/m_{ud}$ and $A=B\ (m_{u}+m_{d})/2=B\ m_{ud}$ where the constant $B$ is determined by the quark condensate $\bar{u}u>$ in the chiral limit, govern the quark mass and quark condensate dependencies of the above one-loop ChPT expressions of the $m_{\pi}^{2},m_{K}^{2},f_{\pi}\ \text{and}\ f_{K}$. The $q$ and $A$ can be expressed below in terms of the masses $m_{\pi},m_{K},m_{\eta}$ and chiral constants $L_{i}$ after inverting the Eqs. (25) and (26) to order $\mathcal{O}(\frac{1}{f^{2}})$ accuracy.
@@
[…]
@@
$\hskip-11.38092pt\frac{1}{f^{2}}\biggl{[}8v_{0}^{(2)}(2m_{K}^{2}+m_{\pi}^{2})(3L_{4}+L_{5})+m_{\pi}^{2}(\mu_{\eta}-3\mu_{\pi})$
$\hskip-11.38092pt-4m_{K}^{2}\mu_{\eta}+\frac{16}{3}\biggl{\{}(6L_{8}-3L_{5}+8L_{7})(m_{\pi}^{2}-m_{K}^{2})^{2}+$
$\hskip-11.38092pt2L_{6}(m_{\pi}^{4}-2m_{K}^{4}+m_{\pi}^{2}m_{K}^{2})+L_{7}(m_{\pi}^{2}+2m_{K}^{2})^{2}\biggr{\}}\biggr{]}\;.$
The $\eta$ mass occurs at the leading order in chiral logarithm i.e. $(m_{\eta}^{(0)})^{2}=(4m_{K}^{2}-m_{\pi}^{2})/3$, hence the functions $f_{\pi}(m_{\pi},m_{K})$, $f_{K}(m_{\pi},m_{K})$ and $M_{\eta}^{2}(m_{\pi},m_{K})$ are applicable only when $4m_{K}^{2}>m_{\pi}^{2}$ . The Eqs.(31) and (32) together with the Eqs.(36), enable the calculation of the QM/QMVT/RQM model couplings $\lambda_{2},c$$m^{2}$ and $\lambda_{1}$ in the $m_{\pi}-m_{K}$ plane.
@@
III Results and Discussions
Several QMVT model studies vac ; lars ; schafwag12 ; chatmoh1 ; guptiw ; vkkr12 ; vkkt13 ; Rai , where the QM model parameters are fixed using the curvature masses of the mesons in the presence of fermionic vacuum fluctuations, have shown that the quark one-loop vacuum fluctuations cause very large smoothing effect on the temperature and chemical potential variations of the non-strange, strange condensates, their derivatives and the mass $m_{\sigma}$ of the scalar $\sigma$ meson. With the parameters obtained using the physical point $\pi,K,\eta\text{ and }\eta^{\prime}$ masses, the 2+1 flavor QMVT and RQM model order parameters, their derivatives and phase diagrams in the $\mu-T$ plane, have been compared extensively in Ref.vkkr22 ; skrvkt24 and it has been shown that the critical end point (where the crossover transition turns second order), shifts to very large chemical potentials and smaller temperatures in the curvature mass parametrized QMVT model where the chiral crossover transition on the temperature axis at $\mu=0$, occurs at a noticeably larger pseudo-critical temperature $T_{c}^{\chi}$ in comparison to the RQM model. Recall that the smoothing effect of the quark one-loop vacuum fluctuation, on the chiral transition is moderate in the RQM model skrvkt24 because on-shell method of renormalizing the parameters in the RQM model, generates a significantly stronger axial $U_{A}(1)$ anomaly ’t Hooft coupling $c$ while the explicit chiral symmetry breaking strength $h_{x}$ ($h_{y}$) in the non-strange (strange) direction becomes weaker by a small amount (a relatively large amount). Note that the $c,h_{x}\text{ and }h_{y}$ do not change in the QMVT model at physical point from their tree level values obtained in the QM model.
@@ III Results and Discussions
Table 2: The scalar $\sigma$ meson mass dependent parameters $\lambda_{10}\ \text{and}\ m_{0}^{2}$ $\{\lambda_{1}\ \text{and}\ m^{2}\}$ of the RQM $\{\text{QMVT}\}$ model, are presented in this table for the $m_{\sigma}=400$ MeV and $m_{\sigma}=530$ MeV.
@@
It is worthwhile to perform the chiral limit study and draw the Columbia plot in the curvature mass parametrized QMVT model where the infrared regularized U(3) ChPT is employed to determine the ($m_{\pi},m_{K}$) dependent numerical values of the $f_{\pi},f_{K}\text{ and }M_{\eta}^{2}$ when one reduces the $m_{\pi}$ and $m_{K}$ while moving away from the physical point. The softening effect of the fermionic vacuum fluctuations in the QMVT model needs to be numerically quantified by calculating its effect on the location of the tricritical point (TCP) at $\mu=0$ in the light chiral limit and the value of the pion (light quark) critical mass $m_{\pi}^{c}$ ($m_{ud}^{c}$) beyond which the chiral transition is a smooth crossover on the three flavor chiral limit line ($m_{\pi}=m_{K}$ i.e. $m_{ud}=m_{s}$ ) in the Columbia plot. In order to see how the chiral limit results are qualitatively and quantitatively influenced by the different methods of treating the quark one-loop vacuum fluctuation when it gets incorporated in the QM Model, the $\mu-m_{K}$, $m_{\pi}-m_{K}$ and the light-strange quark mass $\mu-m_{s}$, $m_{ud}-m_{s}$ planes of the Columbia plots, the $m_{K}^{TCP}(m_{s}^{TCP})$ location of the TCP on the $m_{K}(m_{s})$ axis where the $m_{\pi}(m_{s})=0$ and the $m_{\pi}^{c}$ ($m_{ud}^{c}$) obtained for the QMVT model calculation in the present work will be compared with the corresponding results of the e-MFA:QM model FRG study in the Ref. Resch and the very recent RQM model chiral limit study reported in the Ref. vkt25 .
@@ III.1 QMVT versus RQM model order parameters and its derivatives in the path to chiral limit
The light chiral limit in the QMVT (RQM) model is given by $h_{x}=0$ ($h_{x0}=0$) while the strange chiral limit is given by $h_{y}=0$ ($h_{y0}=0$). The reduced chiral symmetry breaking strengths $h_{x}^{*}$ and $h_{y}^{*}$ ($h_{x0}^{*}$ and $h_{y0}^{*}$) and the corresponding QMVT (RQM) model parameters $\lambda_{2},c,\lambda_{1},m^{2}\ (\lambda_{20},c_{0},\lambda_{10},m_{0}^{2})$ for the chiral limit study, are calculated by using the infrared regularized U(3) ChPT described in the section II.3. The changes in pion, kaon decay constants $f_{\pi}$ and $f_{K}$ are calculated by putting the values of the reduced pion, kaon masses $m_{\pi}^{*}$ and $m_{K}^{*}$ respectively into the Eq. (31) and Eq. (32) while $\text{M}_{\eta}$ is calculated using the Eq.(36). The ratio $\frac{m_{\pi}^{*}}{m_{\pi}}=\frac{m_{K}^{*}}{m_{K}}=\beta$ is kept fixed while the starred $\pi$ and $K$ masses are reduced. The above choice gives us a direct path from the physical point to the chiral limit by just tuning only the fraction $\beta$ instead of randomly changing both the $m_{\pi}$ and $m_{K}$ and ensures that the $\frac{m_{\pi}^{*}}{m_{K}^{*}}$ remains equal to the ratio $\frac{m_{\pi}}{m_{K}}$ Schaefer:09 at the physical point where the ChPT parameter $q=\frac{2\ m_{s}}{m_{u}+m_{d}}=\frac{m_{s}}{m_{ud}}$ as the ratio of the strange $m_{s}$ to the average of light quark mass $m_{ud}$, also gets fixed. The above defined fraction $\beta$, also facilitates the comparison of our results with those of the Refs. Resch ; Schaefer:09 ; vkt25 for the analogous fractions. The reduced light and strange chiral symmetry breaking strengths in the QM/QMVT model are defines as $h_{x}^{*}=(m_{\pi}^{*})^{2}\ f_{\pi}$ and $h_{y}^{*}=\{\sqrt{2}\ f_{K}\ (m_{K}^{*})^{2}-\frac{f_{\pi}}{\sqrt{2}}\ (m_{\pi}^{*})^{2}\}$. The analogous definitions in the RQM model are : $h_{x0}^{*}=(m_{\pi,c}^{*})^{2}\ f_{\pi}$ and $h_{y0}^{*}=\{\sqrt{2}\ f_{K}\ (m_{K,c}^{*})^{2}-\frac{f_{\pi}}{\sqrt{2}}\ (m_{\pi,c}^{*})^{2}\}$ where the reduced pion, kaon curvature masses $m_{\pi,c}^{*}\text{ and }m_{K,c}^{*}$ are found by putting the reduced pion, kaon mass $m_{\pi}^{*}\text{ and }m_{K}^{*}$ values respectively in the expressions Eq. (23) and Eq. (II.2).
@@
The QMVT and RQM model temperature variations of the light and strange condensate $x$ and $y$ have been been compared respectively in the Fig. 1 and Fig. 1 when $\beta=1.0,0.5,0.3642,0.2028\text{ and }0.00725$ for the $m_{\sigma}=530$ MeV. When ChPT expression of $M_{\eta}$ is used as input to determine parameters of the QMVT and RQM models, the $x$ ($y$) plot at the physical point $\beta=1$, overlaps (not shown here), as reported in a very recent RQM model work in the Ref. vkt25 , with the corresponding temperature plot obtained after taking the $M_{\eta}:\text{Expt}$ in input for finding the model parameters whereas the $f_{\pi},f_{K}$ are the same as in the first row of Tables 1 and 2. The light and strange condensate temperature variations in the Fig. 1 and Fig. 1, show a very smooth and delayed melting in the QMVT model for the $\beta=1.0\text{ and }0.5$ while the corresponding melting in the RQM model is sharper and occurs early on the temperature axis. This is evident from the temperature variation of the derivative $\partial x/\partial T$ in the Fig. 2 where the chiral crossover transition in the QMVT model occurs at the pseudo-critical temperature (defined as the peak of the $\partial x/\partial T$ temperature variation) of $T_{c}^{\chi}=162.16$ MeV for $\beta=1$ and $T_{c}^{\chi}=158.32$ MeV for $\beta=0.5$ while a relatively sharper chiral crossover transition, with a higher peak in the $\partial x/\partial T$ temperature plot, occurs in the RQM model at about 23-24 MeV smaller transition temperature of $T_{c}^{\chi}=139.42$ MeV for $\beta=1$ and $T_{c}^{\chi}=133.97$ MeV for $\beta=0.5$ . The $\partial x/\partial T$ peak height in the QMVT model respectively for the $\beta=0.4$ and $\beta=0.3$ in the Fig. 2 turn out to be about five and three times smaller than the height of the $\partial x/\partial T$ peak in the RQM model for the case of $\beta=0.4$ . The reason of the above pattern lies in the fact that the $\beta=0.4$ is close to the RQM model $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{ RQM})}=0.3642\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(50.26,~180.64)$ MeV where the peak in the $\partial x/\partial T$ temperature plot of the Fig. 2, diverges at the $T_{c}^{\chi}=129.33$ MeV and the crossover transition ends at the second order $Z_{2}$ critical point to become first order transition for the $\beta<\beta_{cep}$. The RQM model $\partial x/\partial T$ plots for the $\beta=0.3748$ and $\beta=0.3816$ close to the $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{ RQM})}=0.3642$ are also shown in the Fig. 2 to have a perspective of the divergence. The fraction $\beta$ needs significant reduction to get the $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{QMVT})}=0.2028\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(27.99,~100.59)$ MeV in the QMVT model when the peak height of the $\partial x/\partial T$ plot diverges at larger $T_{c}^{\chi}=145.06$ MeV in the Fig. 2 where the adjacent QMVT model $\partial x/\partial T$ plots are shown for the $\beta=0.2131$ and $\beta=0.2200$ also. The Fig. 1 and 1 show the light and strange order parameter temperature variations also for the $\beta_{cep}$ and $\beta=0.00725$ where the order parameters have a very large first order gap when one reaches very close to the chiral limit. Note that for the smaller $\beta<\beta_{cep}$, the difference in the QMVT and RQM model critical temperatures for the chiral transition decreases from about 24 MeV for the $\beta=1.0\text{ and }0.5$ to 16.35 MeV for $\beta=0.15$ and 13.35 MeV for $\beta=0.00725$. In the exact chiral limit $f_{\pi}=f_{K}$, the $m_{\eta}=0$ and the heavy $m_{\eta^{\prime}}\equiv 825$ MeV signifies the $U_{A}(1)$ anomaly for both the QMVT model and RQM model. The problem of the loss of the spontaneous chiral symmetry breaking ( SCSB) when the $m_{\pi},m_{K}\rightarrow 0$ in the chiral limit for the moderate $m_{\sigma}=400-800$ MeV, gets cured after using the infrared regularized U(3) ChPT scaling Herpay:05 of the $m_{\pi},m_{K},M_{\eta}^{2}$.